Nobeyama Radio Interferometer (NoRI) Analysis Manual ver. 0.0

T. Yokoyama

July 7, 2000

1 Introduction

This is a manual for analysis of the data obtained by Nobeyama Radio Interferometer (NoRI; stop operation in Jul 1992; Nakajima et al. 1980)

For any questions and requests, please send an e-mail to

     nsro-service@solar.nro.nao.ac.jp

The latest information on NoRS is updated on NSRO Web site. The URL is

      http://solar.nro.nao.ac.jp/nori/

2 How to setup

(1) Installation of the SolarSoftware.

Install the SolarSoftware (SSW). If this was not installed, please contact your system manager. The primary distribution site for SolarSoftware is: http://www.lmsal.com/solarsoft/

(2) Setup of your personal environment Include the followings in your environment setup file (~/.cshrc). setenv SSW SSW-

directory^{1}

      setenv SSW INSTR ``nori''
      setenv NORS ${SSW}/radio/nori
      source ${SSW}/gen/setup/setup.ssw
      source ${NORI}/setup/setup.nori

Note that the enviroment variable ${SSW} can be different (Ask your system manager)^{2}. In case you analyze many kinds of data at the same time, define SSW INSTR as follows, e.g.

      setenv SSW INSTR ``nori sxt nors''

3 Where is the Data ?

3.1 Raw Data

Some raw NoRI data are put in the NSRO archive. You can obtain them by anonymous FTP.

1/sgi1/ssw in NSRO

2In NSRO, it is /sgi1/ssw

4 Analysis

4.1 Start Analysis

The NoRI IDL procedures described in this section all depend on the SolarSoftware (SSW). When you start the analysis, start up the SSW/IDL as follows:

unix% sswidl <CR>

5 Read the Data

In order to read the data into IDL session, give the file name as follows:

IDL> file=’./if911202’

IDL> nori rd dat,file,index,data <CR> In order to read the data in specified duration:

IDL> timerange=[’1991-12-2 2:00’,’1991-12-2 4:00’] <CR> IDL> nori rd dat,file,index,data,timerange=timerange <CR>

6 Plot

For plotting, IDL> nori plot,index(0),data(*,0,0) <CR>

A NoRS data archive: anonymous FTP

The Nobeyama Solar Radio Observatory (NSRO/NAOJ) has prepared the anonymous FTP server for the NoRS data archive. The URL is

   ftp://nsro-archive.nro.nao.ac.jp/pub/nori

If you need to transfer a large amount of data , please contact

   nsro-service@solar.nro.nao.ac.jp

We will make and mail a CD-ROM for you. The network capability in Nobeyama is limited.

B Acknowledgement

We thank the following people. SolarSoftware is a software package for analysis of Solar Data on IDL. That is supported under various NASA contracts (SXT/EIT/MDI/TRACE/YPOP etc.) and is all in the public domain. IDL is a product by Research System Inc.

References

Nakajima et al., ”A New 17-GHz Solar Radio Interferometer at Nobeyama”, 1980, PASJ, 32, 639